OpenStudy (anonymous):

If dy/dx = sin^2( piy/4) and y = 1 when x = 0, then find the value of x when y = 3.

OpenStudy (freckles):

do you know how to solve the given differential equation?

OpenStudy (anonymous):

im not too sure how to manipulate dy/dx

OpenStudy (anonymous):

i know i have to isolate the variables and integrate and solve for y right?

OpenStudy (freckles):

\[\frac{dy}{dx}=\sin^2(\frac{\pi}{4} y) \\ \frac{dy}{\sin^2(\frac{\pi}{4} y)}=dx \\ \] like do you know how to ingrate both sides of this equation?

OpenStudy (anonymous):

well the right side is just x but im not to sure on integrating the left side

OpenStudy (freckles):

do you know the derivative of cot(x)?

OpenStudy (anonymous):

-csc^2(x)?

OpenStudy (freckles):

\[\frac{d}{dy} \cot(u)=-\csc^2(u) \cdot \frac{du}{dy} \\ \text{ where } u=u(y)\]

OpenStudy (freckles):

\[\int\limits_{}^{} \csc^2(\frac{\pi}{4} y) dy \\ u=\frac{\pi}{4} y \\ du=\frac{\pi}{4} dy \\ \int\limits \csc^2(u) \frac{4}{\pi} du\]

OpenStudy (freckles):

do you see how to continue

OpenStudy (anonymous):

why did you inverse the pi/4

OpenStudy (freckles):

because if du=pi/4 dy then dy=4/pi du

OpenStudy (freckles):

i just multiplied 4/pi on both sides to isolate dy

OpenStudy (anonymous):

okay okay hold up slow down , why are we integrating csc^2(piy/4)

OpenStudy (freckles):

because that was the left hand side of your equation once you separated the variables

OpenStudy (anonymous):

okay okay i see, give me a second to get my thoughts together

OpenStudy (freckles):

\[\frac{1}{\sin(x)}=\csc(x)\]

OpenStudy (anonymous):

yea but the sine was squared wouldn't that affect that

OpenStudy (freckles):

so you know that equation I just mentioned holds don't you think the equation will still hold if you square both sides

OpenStudy (freckles):

\[(\frac{1}{\sin(x)})^2=\csc^2(x) \\ \frac{1}{\sin^2(x)}=\csc^2(x)\]

OpenStudy (anonymous):

okay sorry i was thinking the square would be negative because it was in the denominator

OpenStudy (anonymous):

i see what you mean though

OpenStudy (anonymous):

okay so we are integrating \[\csc^2(\frac{ \pi y }{ 4 })\]

OpenStudy (anonymous):

oh you were doing substitution , i see

OpenStudy (anonymous):

okay so back to \[\int\limits\limits_{}^{} \csc^2(\frac{\pi}{4} y) dy \\ u=\frac{\pi}{4} y \\ du=\frac{\pi}{4} dy \\ \int\limits\limits \csc^2(u) \frac{4}{\pi} du \] should be do another substitution?

OpenStudy (freckles):

on what? that is just -4/pi*cot(u)+C where u=pi/4y

OpenStudy (anonymous):

\[\frac{ 4 }{ \pi } \int\limits (-\ln(\csc(x) + \cot(x)) + C)^2\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

minus the integral sign

OpenStudy (freckles):

where did you get that answer

OpenStudy (freckles):

remember you yourself said the d/dx cot(x) =-csc^2(x)

OpenStudy (freckles):

or you could say -d/dx cot(x)=csc^2(x)

OpenStudy (anonymous):

are these wrong then?

OpenStudy (freckles):

\[\int\limits\limits\limits_{}^{} \csc^2(\frac{\pi}{4} y) dy \\ u=\frac{\pi}{4} y \\ du=\frac{\pi}{4} dy \\ \int\limits\limits\limits \csc^2(u) \frac{4}{\pi} du \\ \frac{-4}{\pi} \cot(u)+C \\ \frac{-4}{\pi} \cot(\frac{\pi}{4}y)+C\]

OpenStudy (freckles):

nope those all look great

OpenStudy (anonymous):

arghh i feel stupid

OpenStudy (freckles):

anyways we also have to integrate the other side

OpenStudy (freckles):

\[\frac{-4}{\pi} \cot(\frac{\pi}{4} y)+C=x\]

OpenStudy (anonymous):

-_- i was just typing that, okay so now just isolate y right

OpenStudy (freckles):

your second job is to find C for the point (x=0,y=1)

OpenStudy (freckles):

i would not solve for y i think that would cause more work then needed

OpenStudy (freckles):

besides you want to find x

OpenStudy (freckles):

for when y=3

OpenStudy (freckles):

which is the third part of the problem third and final

OpenStudy (anonymous):

\[C = \pi\]

OpenStudy (anonymous):

- pi

OpenStudy (freckles):

\[\frac{-4}{\pi} \cot(\frac{\pi}{4} y)+C=x \\ (0,1)=(x,y) \\ \frac{-4}{\pi}\cot(\frac{\pi}{4}(1))+C=0 \\ \frac{-4}{\pi} \cot(\frac{\pi}{4})+C=0 \\ \frac{-4}{\pi}(1)+C=0 \\ \frac{-4}{\pi}+C=0\]

OpenStudy (anonymous):

:(

OpenStudy (freckles):

it's okay... no frowns here!

OpenStudy (anonymous):

alright well c = 4

OpenStudy (freckles):

\[\frac{-4}{\pi} \cot(\frac{\pi}{4} y)+C=x \\ \frac{-4 }{\pi} \cot(\frac{\pi}{4} y)+\frac{4}{\pi}=x\] last step find x for when y is 3

OpenStudy (freckles):

well no

OpenStudy (freckles):

c=4/pi

OpenStudy (anonymous):

okay let me just take a breath , my frustration is getting the better of me here

OpenStudy (anonymous):

okay so now we just plug c into the solution and plug y = 3 and find x right

OpenStudy (freckles):

right

OpenStudy (freckles):

i will let you do that and I will come back in check in like 5 minutes or less

OpenStudy (anonymous):

okay im getting 8/pi

OpenStudy (freckles):

that sounds right

OpenStudy (freckles):

4/pi+4/pi 2*4/pi 8/pi

OpenStudy (anonymous):

okay great , thanks for the help sorry for being such a dummy :P

OpenStudy (freckles):

i'm going to go for tonight after this problem sleep time but before i go do you have any last questions on this problem?

OpenStudy (freckles):

you aren't a dummy probably just frustrated and trying to learn a new subject

OpenStudy (anonymous):

no im goo thanks again !

OpenStudy (freckles):

the one part is hard enough being frustrated makes it tad harder

OpenStudy (freckles):

anyways peace and i gave you a medal to award you for your effort

OpenStudy (anonymous):

:) night